### v1

# Chirality Amplification and Deracemization in Liquid Crystals

Abstract
Liquid crystal mesophases of achiral molecules are normally achiral, yet in a few materials they spontaneously deracemize and form right- and left-handed chiral domains. One mechanism that drives deracemization is molecular shape fluctuations between axial chiral conformations, where molecular interactions favor matching chirality and promote helical twist. Cooperative chiral ordering may also play a role in chirality amplification, as when a tiny fraction of chiral dopant drives a nematic phase to become cholesteric. We present a model of cooperative chiral ordering in liquid crystals using Maier-Saupe theory, and predict a phase diagram with a deracemized cholesteric phase as well as racemic nematic and isotropic phases. Our model also predicts chirality amplification in the nematic phase, which may be observed even in materials where the deracemization transition is preempted by a transition to another phase. We compare these results with Monte Carlo simulation studies of the switchable chiral Lebwohl-Lasher model, where each spin switches between right- and left-handed chiral states. Simulation results validate the predicted phase diagram, demonstrate chiral amplification in the racemic nematic phase, and reveal coarsening dynamics in the deracemized phase. Our results suggest that cooperative chiral ordering via molecular shape transitions is a common mechanism in liquid crystals.
Chiral symmetry breaking and chirality amplification are attracting considerable interest in biology, chemistry, materials science and physics due to their relevance to the origin of life and single-handedness of molecules of biological origin [1]. In both, chirality at the molecular scale can be translated to the supramolecular and even the macroscopic scale [2]. For instance, a very small enantiomeric excess of chiral molecular units produces helical polymers of definite handedness, not only in covalent polymers [3] but also in supramolecular ones [4, 5]. Mixing tiny fractions of a chiral compound with achiral molecular building blocks does the same. The former is called the “majority rules” principle and the latter the “sergeants-and-soldiers” effect [6]. Likewise, a small excess of one of the enantiomers or the addition of very small amounts of a chiral compound referred to as “chiral doping” is sufficient to break chiral symmetry in liquid crystals [7, 8]. Recently, lab-synthesized bacteria with biochemical components of opposite chirality to those observed in nature, so-called “mirror cells”, have recently been identified as a potentially dangerous risk to life on Earth [9].
Generally, chirality amplification is thought to arise via co-operative ordering. In polymers, cooperative chiral ordering can occur via long-range correlations between left- and right-handed helically bonded states of neighboring molecular units along the backbone of a polymer chain [4, 10]. These long-range correlations arise because of an underlying Ising-like phase transition that in (quasi) one-dimensional systems is suppressed by thermal fluctuations. We argue here that via a similar mechanism, chiral interactions cause molecular conformations of nearby mesogens in a liquid crystalline fluid to become correlated. This cooperative chiral ordering of molecules gives rise to a diverging chiral susceptibility and must be at the root of chiral symmetry breaking in nematic liquid crystals. We show here that the helical twisting power associated with adding a chiral dopant to a nematic must be inversely proportional to the distance to a second-order phase transition, in which dynamic left-handed and right-handed conformers phase deracemize into cholesteric liquid crystals with opposite handedness. Our calculations show that the helical twisting power may exhibit a complex temperature dependence, dependent on the properties of both the mesogens and the dopants. Our work also explains why some liquid crystalline materials consisting of non-chiral mesogens spontaneously deracemize and form right- and left-handed chiral domains [11], noting also that pretransitional fluctuations signaling deracemization have been seen in simulations of achiral bent-core model mesogens [12]. We are not aware of any experimental work showing that deracemization has yet been observed in more conventional nematic liquid crystals, but argue that the chiral transition that we identify may be hidden below the transition temperature to a smectic or crystalline phase.
We first discuss the main ingredients and findings of our mean-field theory, and next discuss those of our Monte Carlo studies. We end with a short discussion and conclusions, and suggestions for further study.
Results
Maier-Saupe Theory
As a starting point we take the Maier-Saupe theory of van der Meer and collaborators [13], and extend it to mixtures of $\pm$ enantiomers: compounds with axial chiralities P and M that prefer opposite cholesteric wave numbers. The fraction of $\pm$ components we denote $f_{+}=1-f_{-}$, a quantity that we shall treat as an equilibrium variable. We note in this context that, for example, in non-chiral biphenyl mesogens these enantiomers may inter convert dynamically [7] with estimated barriers between 3 and 5 times the thermal energy (at room temperature) [14]. Interactions between the $+$ components favor a dimensionless helical wave number $+q$ and those between the -- components a helical wave number of $-q$. In this model, interactions between $+$ and -- components do not produce a helical pitch. For simplicity we presume all nematic interactions to be the same between the various components and described by a single coupling constant $J>0$. In contrast to the earlier work of Terenteva and Shiyanovskii [15], our model explicitly deals with the chiral interaction between the mesogens and predicts a cholesteric pitch. Our Monte Carlo simulations, discussed in more detail in the next section, are based on the same type of interaction potential [13].
Accounting for the mixing entropy of the two species of compound, we obtain for the Helmholtz free energy $F$ per particle
$\begin{split}F&=k_{\mathrm{B}}T\langle\ln P\rangle-\frac{1}{2}JS^{2}\\ &+\frac{1}{2}\left(Jq^{2}-Kq\right)S^{2}f_{+}^{2}+\frac{1}{2}\left(Jq^{2}+Kq\right)S^{2}f_{-}^{2}\\ &+k_{\mathrm{B}}Tf_{+}\ln f_{+}+k_{\mathrm{B}}Tf_{-}\ln f_{-}-2gf_{+},\end{split}$ (1)
[…]
The third and fourth terms in Eq. 1 arise due to the chiral interaction that induces a helical deformation of the director field with dimensionless wave number $q$. The quadratic terms in $q$ account for the reduction of the nematic interaction due to the twisting of the director field, whilst the terms linear in $q$ measure how strongly the chiral interaction is favored by the as yet unknown degree of helicity of the director field. These terms follow from a mean-field type of analysis of a chiral Lebwohl-Lasher type of interaction potential between model mesogens [13]; see also Eq. 5.
The last three terms describe the mixing entropy of the species $\pm$, and a term that breaks chiral symmetry with $g$, a free energy difference between the two enantiomers. The factor of 2 we introduce to connect with our simulations in which the equivalent term is $g\eta\propto 2gf_{+}$. This chiral biasing term may, e.g., represent a chemical potential difference that in a grand canonical ensemble regulates the enantiomeric excess $\eta=2f_{+}-1=1-2f_{-}$. Alternatively, it may represent a free energy difference between the two enantiomers, induced by the action of circularly polarized light or the binding of a chiral dopant [7]. In the latter case we expect that $g$ is proportional to the concentration of dopant, the consequences of which we investigate in more detail below.
The optimal orientational distribution function $P(\cos\theta)$ functionally minimizes the free energy, and the equilibrium values of the chiral wave vector $q$ and the fraction $+$ and -- states $f_{+}=1-f_{-}$ also minimize the free energy. For $g=0$ and a sufficiently weak chiral interaction, the transition between a racemic isotropic and a racemic nematic phase occurs at the clearing temperature $T_{\mathrm{IN}}$ for which $J/k_{\mathrm{B}}T_{\mathrm{IN}}=4.54$ where $S=0.43$ in the nematic phase [16], when the free energies of the coexisting isotropic and nematic phase are equal. Let us consider the fluid to be in the nematic phase with temperature $T<T_{\mathrm{IN}}$ and order parameter $S>0.43$.
If we minimize the free energy Eq. (1) with respect to $q$, we obtain
$x\equiv\frac{q}{q_{0}}=\left(\frac{f_{+}^{2}-f_{-}^{2}}{f_{+}^{2}+f_{-}^{2}}\right),$ (2)
with $q_{0}=K/2J$ the intrinsic cholesteric wave number of the enantiomers and $x$ the ratio of the equilibrium value of $q$ and that of the intrinsic one, $q_{0}$. We require that $|q_{0}|\lesssim 1$ for the pitch not to drop below a microscopic length, which corresponds to $|K|/J\lesssim 2$. As expected, we find for racemic mixtures, with $f_{+}=f_{-}=1/2$ and enantiomeric excess $\eta=0$, $x=0$ and hence $q=0$. If on the other hand $f_{+}=1-f_{-}=0,1$ and $\eta=\pm 1$, we retrieve $x=\pm 1$ and $q=q_{0}$, which is the equilibrium value of the cholesteric wave number of the enantiomerically pure material within our model.
We also need to minimize the free energy Eq. 1 with respect to $f_{+}=1-f_{-}$, to find
$\begin{split}\chi\left(\frac{1}{2}x^{2}-x\right)f_{+}-\chi\left(\frac{1}{2}x^{2}+x\right)f_{-}\\ +\ln\left(\frac{f_{+}}{f_{-}}\right)-\frac{2g}{k_{\mathrm{B}}T}=0,\end{split}$ (3)
where the strength of the chiral interaction is given by $\chi\equiv(K/J)^{2}S^{2}J/2k_{\mathrm{B}}T=2.27\hskip 2.84544pt(T_{\mathrm{IN}}/T)\hskip 2.84544pt(K/J)^{2}S^{2}$. In the absence of a chiral interaction between the mesogens, $\chi=0$. This may be the case if $K=0$ and $q_{0}=0$, or if the fluid is in the isotropic phase where $S=0$. In both cases, the fractions $+$ and -- enantiomers obey a simple Boltzmann distribution with $f_{+}/f_{-}=\exp(2g/k_{\mathrm{B}}T)$. The corresponding enantiomeric excess $\eta=\tanh g/k_{\mathrm{B}}T$ behaves like the average spin state of the Ising model of ferromagnetism, where $g$ takes on the role of magnetic field. In our model, the spins do not directly couple, only indirectly through the chiral interaction between the species of nematogen.
Figure 1: Left: Enantiomeric excess $\eta$ (blue – inner curve) and ratio $x$ of the cholesteric wave number $q$ and the maximal value $q_{0}$ (red – outer curve) in the co-existing deracemized cholesteric phases as a function of the dimensionless magnitude of the chiral interaction between the nematogens $\chi$ defined in the main text. Right: phase diagram as a function of the temperature $T/T_{\mathrm{IN}}$ scaled to the nematic transition temperature $T_{\mathrm{IN}}$ and the ratio of the chiral and nematic interaction strengths $|K|/J$. The vertical line in red indicates the isotropic-nematic transition. The curved line in blue demarcates the chiral transition between the deracemized cholesteric phase and the racemic nematic phase for $T<T_{\mathrm{IN}}$ and the racemic isotropic phase for $T>T_{\mathrm{IN}}$, which cross at the tricritical point (indicated).
For $g=0$, there is no overall enantiomeric excess: the mesogens have no inherent preference for either enantiomeric state. However, even if $g=0$ local chiral symmetry may still be broken. This happens if $\chi>1$, as is shown in Fig. 1 (left) where we present a numerical solution of Eqs. 2 and 3 for $g=0$. For $\chi\leq 1$ the enantiomeric excess $\eta\equiv 2f_{+}-1$ and cholesteric wave number $x=q/q_{0}$ are both equal to zero. This represents the achiral, racemic nematic phase. For $\chi\geq 1$, local chiral symmetry is broken and the racemic nematic phase deracemizes into two co-existing cholesteric phases with opposite cholesteric wave number ±$\pm|x|$ and opposite enantiomeric excess ±$\pm|\eta|$. In our model, the chiral or deracemization transition is continuous, and is reminiscent of the spontaneous magnetization of the Ising model below the Curie temperature.
The condition $\chi_{\mathrm{c}}=(K/J)^{2}S^{2}(T_{\mathrm{c}})J/2k_{\mathrm{B}}T_{\mathrm{c}}=1$ identifies a critical chiral transition or deracemization temperature $T_{\mathrm{c}}/T_{\mathrm{IN}}=2.27\hskip 2.84544pt(K/J)^{2}S^{2}(T_{\mathrm{c}})$. For weak chiral interaction, so small values of $|K|/J$, $T_{\mathrm{c}}<T_{\mathrm{IN}}$. If $T_{\mathrm{c}}\ll T_{\mathrm{IN}}$ the chiral transition might be obscured by a transition to, say, a smectic or a crystalline phase. On the other hand, $T_{\mathrm{c}}$ moves towards $T_{\mathrm{IN}}$ with increasing chiral interaction $|K|$ as is shown in our phase diagram Fig. 1 (right). The deracemization transition crosses the nematic transition at a tricritical point for $|K|/J=1.54$. For $|K|/J>1.54$ there is a direct transition from a racemic isotropic phase to two deracemized cholesteric phases of equal volume. The phase diagram we obtain from our simulations, shown in Fig. 3 (top), is in good agreement with that from our mean-field theory of Fig. 1 (right).
Not surprisingly, we obtain mean-field critical exponents near the chiral transition. In fact, for $T\rightarrow T_{\mathrm{c}}$, Eqs. 2 and 3 can be solved analytically giving $x\sim\pm 4\hskip 2.84544pt(3/28)^{1/2}(\chi-1)^{1/2}\pm\propto(T_{\mathrm{c}}-T)^{1/2}$ and $\eta\sim\pm 2\hskip 2.84544pt(3/28)^{1/2}(\chi-1)^{1/2}\propto\pm(T_{\mathrm{c}}-T)^{1/2}$ for $\chi\geq 1$ or $T\leq T_{\mathrm{c}}$, and $\eta=x=0$ for $\chi\leq 1$ or $T>T_{\mathrm{c}}$. Arguably, the existence of a second order chiral transition explains the phenomenon of chirality amplification because near a critical point susceptibilities tend to become large and in fact diverge at the critical point. This should show up in the response of the chiral order parameters $x$ and $\eta$ to the chiral biasing field $g$.
This point is illustrated in Fig. 2 (a). The figure shows our solution to Eqs. 2 and 3 for the enantiomeric excess $\eta$ as a function of $g/k_{\mathrm{B}}T$ for different scaled temperatures $T/T_{\mathrm{IN}}$ and fixed chiral interaction $K/J=0.333$. Recall that the chiral interaction strength $\chi$ can be written as $\chi=\frac{1}{2}(K/J)^{2}S^{2}(J/k_{\mathrm{B}}T)$ where $J/k_{\mathrm{B}}T=4.54\hskip 2.84544ptT_{\mathrm{IN}}/T$. This means that the enantiomeric excess $\eta$ is not only a function of the dimensionless biasing field $g/k_{\mathrm{B}}T$, but also a function of $K/J$ as well as of $T/T_{\mathrm{IN}}$ because of the nematic order parameter $S=S(T/T_{\mathrm{IN}})$ also depends on the temperature. The values of $S(T/T_{\mathrm{IN}})$ we obtain from a numerical solution of the self-consistent field equation of Maier-Saupe theory. For the temperatures $T/T_{\mathrm{IN}}$ shown, $\chi$ is smaller than unity and the nematic remains racemic. Still, we do see a significant increase in the enantiomeric excess relative to the hyperbolic tangent relation describing the case $K=0$, and more so the lower the temperature. The trends of our theoretical predictions agree with those from our simulations, as Fig. 2 (b) confirms. See also the Supporting Information, Fig. S1, in which we investigate the impact of larger and smaller values of $K/J$. For $\chi>1$, the Maier-Saupe theory predicts hysteresis of the enantiomeric excess $\eta$ as a function of the biasing field $g$, with metastable positive (negative) valued $\eta$ for negative (positive) values of $g$ (not shown).
Figure 2: Enantiomeric excess $\eta$ for a fixed strength of the chiral interaction $K/J=0.333$, as a function of the dimensionless chiral biasing potential $g/k_{\mathrm{B}}T$ for different temperatures $T/T_{\mathrm{IN}}$. (a) Predictions from Maier-Saupe theory. The corresponding values of the chiral interaction strength $\chi$ are $0.825,0.634,0.412$ and $0.239$ from top to bottom. (b) Results of Monte Carlo simulations. Indicated is also the ideal hyperbolic tangent behavior expected in the absence of a chiral interaction and $K=0$.
In the limit $\chi\rightarrow 1$ the susceptibilities of the enantiomeric excess $\chi_{\eta}\equiv\lim_{g\rightarrow 0}\partial\eta/\partial g$ and that of the cholesteric wave number $\chi_{x}\equiv\lim_{g\rightarrow 0}\partial x/\partial g$ both actually diverge. From Eqs. 2 and 3, we find that for $\chi\uparrow 1$, $\chi_{\eta}=\frac{1}{2}\chi_{x}=\frac{1}{4}(1-\chi)^{-1}/k_{\mathrm{B}}T$, whilst for $\chi\downarrow 1$, $\chi_{\eta}=\frac{1}{2}\chi_{x}\sim\frac{7}{24}(\chi-1)^{-1}/k_{\mathrm{B}}T$. So, in both cases we find $\chi_{\eta}\propto|T-T_{\mathrm{c}}|^{-1}$. Our simulations produce the same mean-field exponent for $\chi_{\eta}$ (Fig. S4) albeit that we have not performed any finite-size analysis to obtain an accurate value. Finally, accounting in the free energy for square gradient terms in the variables $x$ and $f_{+}=(\eta+1)/2$ produces two correlation lengths proportional to $1/\sqrt{1-\chi}$ confirming that chiral conformational states are indeed “contagious” and passed on to neighboring mesogens through the chiral interaction.
The susceptibility $\chi_{x}$ associated with the cholesteric wave number may be interpreted to represent what is known as the helical twisting power of a chiral dopant [17]. A tiny induced preference for a particular enantiomeric form of the nematogens becomes strongly enhanced by the coupling to the chiral symmetry breaking transition. This may well explain why so little chiral dopant is needed to turn a nematic into a cholesteric [7, 8]. That different dopants produce varying helical twisting powers if added to the same mesogen may be understood in terms of unequal surface interactions with the director field [18], expressed in our model in different values of the binding free energy, $g$. In fact, for large enough binding energies, $g$ may be replaced by $gX$ with $X\ll 1$ the mole fraction of dopant, in which case the helical twisting power, $\beta$, may be written as
$\beta=\frac{1}{1-\chi}\left(\frac{2g}{k_{B}T}\right)\frac{q_{0}}{2\pi}.$ (4)
The helical twisting power efficacy of the same dopant in different nematogens may be interpreted in terms of the distance in temperature to the chiral transition of that nematogen, yielding different values of $\chi$, and differences in the binding free energy, $g$, and the intrinsic helical wave number, $q_{0}$. Changes in the binding free energy between different dopants result in different helical twisting powers dependent on the nematogen. Novel in this expression for the helical twisting power is the pretransitional enhancement factor $(1-\chi)^{-1}$ not accounted for in other approaches [18].
It turns out to be nontrivial to evaluate the temperature dependence of the helical twisting power, as $\chi$, $g$ and $q_{0}$ all depend on the temperature. For our model, $q_{0}=K/2J$ is often thought to not depend on temperature [13] but this presumes that energies $K$ and $J$ are not free energies. Since the interaction energy is based on a coarse-grained model, $K$ and $J$ must in fact be interpreted as free energies in which case both are in principle temperature dependent. If we ignore this for the sake of argument, and presume $K$ and $J$ to be a constant of temperature, then we deduce from Eq. 4 that $\beta$ should decrease with temperature under most (but not all) conditions, because higher temperatures imply larger distance to the critical deracemization temperature and smaller values of $\chi$. This is often seen experimentally but not always [19]. However, if the binding of the dopant to the nematogen is entropy- rather than enthalpy-driven, and $g/k_{B}T$ increases with temperature, then it may actually increase with temperature. This could explain recent experimental observations [20]. We refer to the Supporting Information for a more detailed discussion of the factors that influence the temperature dependence of the helical twisting power $\beta$.
Simulation Results
To explore cooperative chiral ordering and spontaneous deracemization, we study by means of Monte Carlo simulation a switchable chiral Lebwohl-Lasher model as described by the following Hamiltonian:
$H=\sum_{\langle i,j\rangle}\left[-J\hskip 1.42271ptP_{2}\left(\hat{u}_{i}\cdot\hat{u}_{j}\right)\right.\\ \left.-\frac{1}{2}K\left[(\hat{u}_{i}\times\hat{u}_{j})\cdot\hat{r}_{ij}\right](\hat{u}_{i}\cdot\hat{u}_{j})(\eta_{i}+\eta_{j})-g\eta_{i}\right],$ (5)
where the sum is over neighboring pairs $\langle i,j\rangle$ of lattice sites, $J$ and $K$ have the same meaning as before, as does the biasing field $g$, and $\hat{r}_{ij}$ denotes the unit vector connecting these two sites. Associated with each site are two order parameters: the unit orientation vector $\hat{u}_{i}$ and the chirality $\eta_{i}=\pm 1$ describing the left or right handedness of the mesogens. The first term in Eq. 5 represents the usual Lebwohl-Lasher potential [21, 22], and the second term is a chiral interaction term recently explored in a different context by Elsasser and Kuhnhold [23] without the local switchable chirality that is present in our model. The coupling term $\eta_{i}+\eta_{j}$ promotes twist between neighboring “spins” for pairs of like chirality $\eta_{i}=\eta_{j}$ but not between neighboring spins of opposite chirality $\eta_{i}\neq\eta_{j}$. More simulation details are available in the methods section. Simulations were performed on the National Center for Supercomputing Applications’ Delta cluster through the NSF ACCESS program [24]. A copy of the simulation code is available at [25].
Figure 3: Phase diagrams as a function of the chiral interaction strength $K/J$ and scaled temperature $T/T_{\mathrm{IN}}$ obtained by means of Monte Carlo simulations. (a) Heat map of the logarithm of the specific heat, $\mathrm{log}_{\mathrm{10}}C_{V}$, and (b) that of the chiral susceptibility $\chi_{\eta}$. The latter cannot distinguish between the racemic isotropic and nematic phases.
We simulate the model using an $N\times N\times N$ cubic lattice with $N=128$. To explore the phase diagram, the chiral interaction strength, $K/J$, was varied from $K/J=0.0$ to $K/J=2.0$ with a total of $13$ values. For each value of $K/J$, the temperature was varied from $T/T_{IN}=2.0$ to $T/T_{IN}=0.067$ with $2\times 10^{6}$ Monte Carlo steps at each of the $30$ temperature values. For each set of parameters, we used the last $6.25\%$ of configurations to calculate specific heat and chiral susceptibility, with configurations separated by 100 Monte Carlo steps per spin. The specific heat was calculated from the energy fluctuations as $C_{V}=\left(\langle E^{2}\rangle-\langle E\rangle^{2}\right)/k_{\mathrm{B}}T^{2}$. The chiral susceptibility is calculated from the fluctuations in the enantiomeric excess $\eta$, $\chi_{\eta}=N^{3}\left(\langle\eta^{2}\rangle-\langle\eta\rangle^{2}\right)/k_{\mathrm{B}}T$. The enantiomeric excess was calculated as the average chirality per lattice site, in direct analogy to the average spin state in the Ising model.
The phase diagram obtained from our simulations is shown in Fig.3 (a), showing peaks in the heat capacity that indicate transitions between racemic isotropic, racemic nematic and deracemized cholesteric phases, in close agreement with predictions of our Maier-Saupe theory of Fig. 1. We find a racemic isotropic phase at high temperature, a racemic nematic phase at low temperature and sufficiently weak chiral interaction $K/J$, and a spontaneously deracemized cholesteric phase at low temperature and high $K/J$. The intersection of the phase boundaries suggests the existence of a tricritical point, also predicted by our Maier-Saupe theory. We in addition plot the chiral susceptibility in Fig. 3 (b), which can only show the phase boundary of the spontaneously deracemized cholesteric phase. It confirms that near the transition, chiral fluctuations become large and hence also the chiral susceptibility.
Even not all that close to the transition, the increased susceptibility can be quite significant as Fig. 2 (b) confirms. In the figure, the enantiomeric excess $\eta$ is shown as a function of the dimensionless chiral bias field for different temperatures but fixed chiral interaction strength $K/J=0.333$. In our Monte Carlo simulations, we varied $g/k_{\mathrm{B}}T$ from $0.0$ to $0.25$, $K/J$ from 00 to $0.5$, and $T/T_{IN}$ from $0.668$ to $0.267$, where we annealed $2\times 10^{6}$ Monte Carlo steps at each temperature. In agreement with our theory, we find that the chiral susceptibility goes up with decreasing temperature and with increasing chiral interaction strength $K/J$, as the Fig. S1 shows. According to our theory, both $T$ and $K/J$ enter because $\eta$ not only depends on $g/k_{\mathrm{B}}T$ but also on the value of $\chi$ that itself is a function of $T/T_{\mathrm{IN}}$ and $K/J$.
Figure 4: (a-b): Simulation results showing local orientation and chirality configurations at fixed temperature $T/T_{IN}=0.067$ and varied values of chiral interaction, $K/J$. A horizontal $XZ$ cross-section is shown with the local unit orientation field $\hat{u}_{i}$ colored blue and red corresponding to $\eta_{i}=\pm 1$. Simulation box size $N=128$. (c): Simulation of size $N=480$ after $2\times 10^{6}$ steps showing alternating chiral domains. (d): Double-logarithmic plot of the average layer thickness, $\langle L(t)\rangle$, as a function of time, $t$, in number of steps for the box size $N=480$. Stripes of alternating chirality where counted across the $101$ plane every $1\times 10^{5}$ steps. The dashed line shows a curve fit of $\langle L\rangle\propto t^{\alpha}$ producing a scaling exponent $\alpha=0.379\pm 0.016$.
What Maier-Saupe theory cannot show, and our simulations can, is the spatial structure of the phases and their coarsening behavior. Fig. 4 (a) and (b) show $xz$ cross-sections of a system with size $N=128$, with open boundary conditions on all sides. Both simulations were annealed to a temperature of $T/T_{IN}=0.067$ starting from $T/T_{IN}=1.0$, in $30$ steps of $2\times 10^{6}$ Monte Carlo steps each. In the figures, right-handed ${\eta_{i}}=+1$ spins are shown in red and left-handed ${\eta_{i}}=-1$ spins are shown in blue. Fig. 4(a) shows the racemic nematic phase at low value of $K/J=0.167$, with short range chiral correlations, but no long range order and no induced cholesteric twist. Fig. 4(b) is at a higher value of $K/J=0.5$, showing the deracemized phase with layers of cholesteric of approximately two pitch lengths each. A more detailed image can be found in Fig. S3 demonstrating that under these conditions the interface between left- and right-handed deracemized cholesteric layer is indeed very sharp. The cholesteric domains form with the pitch axis oriented along a body diagonal of the cubic lattice, see also Fig. 4 (c). This orientation of the pitch axis is allowed by open boundary conditions and is lower in energy than a pitch oriented along the $x$, $y$ or $z$ axes of the cubic simulation volume.
To investigate coarsening behavior in the deracemized phase, we carried out a simulation of a larger system with $N=480$ and open boundary conditions, for a fixed chiral interaction ${K/J}=0.5$ and a temperature $T/T_{IN}=0.067$ that falls in the deracemized cholesteric region of the phase diagram. We followed the coarsening over $2\times 10^{6}$ Monte Carlo steps per spin. Our findings are shown in Fig. 4 (c) and (d), where layers with ${\eta_{i}}=1$ are again shown in red and with ${\eta_{i}}=-1$ in blue. Layers with alternating chiral sense form within the first $1\times 10^{5}$ time steps after starting from an initial random configuration as seen in Fig. S5 (a). These widen with time as can be seen in more detail a in the cross sections shown in Fig. S2 and Fig. S5. Initially, multiple domains form with layer normals aligned along different body diagonal axes. Over time, those domains grow with time, the average layer width increases, and the number of layers drop. Layers annihilate via two mechanisms: (i) via motion and annihilation of edge dislocations in the layer structure, and (ii) via layer annihilation at corners of the cubic box facilitated by the open boundary conditions. The average layer thickness $L$ we find to increase continuously as $L\sim{t^{\alpha}}$ with scaling exponent $\alpha={0.379}\pm{0.016}$ as shown in Fig. 4 (d). Because layer formation and coarsening are driven by an intrinsically anisotropic surface tension, we do not expect this scaling exponent to be universally valid [26, 27].
As Fig. 4 (c) shows, the domain structure is actually quite complex, each with layer normals oriented along different body diagonals. Layers formed in early stages of pattern formation have thicknesses less than the cholesteric pitch, increasing to several pitch lengths during coarsening. See Fig. S3 and Fig. S6 b). Layers less wide than about a pitch do not have a simple, single twisted cholesteric structure but exhibit a double-twist structure. Layers annihilate by shrinking to zero thickness at a corner or edge of the simulation box, or via motion and annihilation of edge dislocations. The system remains in a poly-domain structure after $2\times{10^{6}}$ Monte Carlo steps per spin. The dynamics of micro-structural evolution is shown in Supplementary Video S1 and in Fig. S5. Cross-sectional images are shown in Fig. S2.
Finally, for values of $K/J=0.833$ and larger, a more complex, disordered double twist geometry without a uniform twist axis emerge. The defect-rich micro-structures suggest the formation of blue phases, which have not previously been observed in studies of the chiral Lebwohl-Lasher model. We plan to investigate the details of the blue phase in the uniform chirality Lebwohl-Lasher model in future work.
Discussion
Spontaneous deracemization may occur by a variety of mechanisms [28], and is not restricted to racemic mixtures of chiral molecules. In fact, achiral mesogens can form chiral phases even in the absence of intramolecular degrees of freedom. For instance, bent-core molecules self-assemble into smectic phases with a chiral arrangement of the layer normals, the molecular dipole orientations, and molecular tilts [29, 30]. Also, twist-bend nematic phases form without reliance of intramolecular degrees of freedom [31, 32, 33, 34, 35]. Intramolecular degrees of freedom can also drive deracemization [15]. While molecular chirality typically arises due to the presence of a chiral center, an achiral molecule with degenerate right- and left-handed twist conformations can be described as having axial chirality [8]. Here two mirror-image chiral conformations differ only by the sign of axial twist in the molecular core.
In the model we have presented, our focus is on mesophases of molecules with axial chirality where the energy barrier between mirror-image conformations is sufficiently small relative to the thermal energy, enabling switching between chiral states. Our model also assumes that interactions of neighboring molecules promote agreement in their axial chiral state and predicts segregation into chiral domains. Wilson et al.  have clearly demonstrated this mechanism via atomistic molecular simulations of flexible bent-core mesogens [12]. In related work there is evidence that intramolecular degrees of freedom in bent-core molecules enables them to form chiral conformations with resulting impact on their phase behavior as well as their chiral effect as dopants on other mesophase materials [36].
In summary, by making use of a relatively simple Maier-Saupe theory that allows for a thermodynamic equilibrium to exist between conformers of opposite handedness, we are able to explain chirality amplification in nematic liquid crystals. According to our model, chirality amplification is caused by an underlying tendency of enantiomers with opposite handedness to demix, which makes them susceptible to a chiral field even when the second order demixing transition is obscured by a transition to another liquid or solid crystal phase. In follow-up work, we intend to study the phase behavior of our model in much more detail and focus on potential dynamical effects associated with the chiral hysteresis predicted by our theory, in analogy to the hysteresis observed in the magnetization of ferromagnets.
Methods
Theory
Following van der Meer et al. [13], we find from a mean-field type of analysis of the impact of a chiral Lebwohl-Lasher type of interaction potential between model mesogens to lead to expressions for the self-consistent fields experienced by species $\alpha=\pm$ due to the presence of species $\beta=\pm$ to read
$V_{\alpha,\beta}(\theta)=\left(-J(1-q^{2})-qK_{\alpha,\beta}\right)SP_{2}(\cos\theta)f_{\beta}.$ (6)
The chiral coupling constants $K_{\alpha,\beta}$ are $K_{+,+}=K$ between the $+$ species, $K_{-,-}=-K$ between the -- species and $K_{+,-}=K_{-,+}=0$ between the species with opposite handedness. In Eq. 6, we have absorbed unimportant numerical constants in the definitions of the interaction strengths $J$ and $K$ as well as the cholesteric wave number $q$ that in our description is made dimensionless by multiplication with a microscopic length scale [13]. The free energy of interaction follows from summing all self-consistent fields experienced by species $\alpha$ multiplied with the appropriate fraction $f_{\alpha}$, and accounting for double counting. Adding the Gibbs entropy for the orientational distribution and that for the distribution over the species $\alpha$, as well as the contribution of the biasing field $g$, then produces, after some rearranging, Eq. 1.
Simulation Methods
To explore cooperative chiral ordering and spontaneous deracemization by means of Monte Carlo computer simulation, we study a switchable chiral 3D Lebwohl-Lasher model based on the Hamiltonian of Eq. 5. We perform Monte Carlo simulations of the model via the Metropolis algorithm, where at each Monte Carlo step we separately vary the local orientation $\hat{u}_{i}$ and chiral state parameter $\eta_{i}$. We use a novel method for trial spin reorientation which realizes a smoother reorientation of the spin, rather than the Barker-Watts technique [37] used in [22]. To perform a random walk on a unit sphere, we calculate two vectors orthogonal to the orientation vector $\hat{u}_{i}$ and tangent to the unit sphere, then reorient the spin to a random position on the edge of a disk of radius $d$, and normalize the orientation vector to unit length. The disk size $d$ represents the step size and is adjusted to keep the Monte Carlo acceptance rate in the range of 0.4 to 0.6 during equilibration at each temperature, and after that is held fixed. We initialize each simulation in the racemic isotropic phase.
Instead of periodic boundary conditions, we use open boundary conditions. This allows twist without any constraints imposed by periodic boundary conditions, and does not impose any preferred orientation of the twist axis. We find that the spontaneously chosen twist axis is typically not along the $x,y$ or $z$ axis, as in fact was assumed in [23], but along a body diagonal of the cubic lattice. We also note that open boundary conditions allow topological defects to nucleate and annihilate at the free surfaces.
Lattice-based Monte-Carlo simulations are a computationally “embarrassingly parallelizable” problem [38]. Our Monte Carlo code is highly optimized to run on graphics processing units (GPUs) using a combination of Fortran and OpenACC for the bulk of the computation, and Python for simulation management, data processing, analysis, and visualization. Simulations were performed on the National Center for Supercomputing Applications’ Delta cluster through the NSF ACCESS program [24]. A copy of the simulation code is available at doi.org/10.5281/zenodo.14781846. [25].
Acknowledgements.
This work used the Delta supercomputer at the National Center for Supercomputing Applications through allocation PHY240256 from the Advanced Cyberinfrastructure Coordination Ecosystem: Services & Support (ACCESS) program, which is supported by U.S. National Science Foundation grants #2138259, #2138286, #2138307, #2137603, and #2138296. This research was supported by Grant No. 2022197 from the United States-Israel Binational Science Foundation (BSF). We thank Dr. Mark Wilson (Durham University, UK) and Dr. Daniela Cywiak (University of Guanajuato, Mexico) for stimulating discussions.
Appendix A Supplementary Information
Dopant binding to a dense mesogen fluid
A simple model for the binding of chiral dopants to achiral mesogens may be provided by a Flory-Huggins type of fluid lattice model. The total number of lattice sites equals $M$. Let $\phi$ be the volume fraction of free mesogens, and $\phi_{*}$ and $\phi_{2}$ that of free dopant molecules and dimers of mesogens and dopants. The volume fractions add up to unity, $\phi+\phi_{*}+\phi_{2}=1$. Dimers take up two lattice sites, whilst free mesogens and free dopants each occupy only a single lattice site. The overall volume fraction of free and bound dopants is $\Phi_{*}$ and that of the free and bound mesogens is $\Phi=1-\Phi_{*}$. Hence, we have $\phi_{*}=\Phi_{*}-\frac{1}{2}\phi_{2}$ and $\phi=\Phi-\frac{1}{2}\phi_{2}$.
We ignore any net non-bonded interaction between mesogens, dopants and dimers, implying that the mesogen fluid acts as an ideal solvent to the dopants and dimers. We do account for a binding free energy $-g/k_{\mathrm{B}}T<0$ associated with each dimer. The Helmholtz free energy $F$ of the fluid mixtures in that case reads
$\frac{F}{Mk_{\mathrm{B}}T}=\phi\ln\phi+\phi_{*}\ln\phi_{*}+\frac{1}{2}\phi_{2}\ln\frac{1}{2}\phi_{2}-\frac{1}{2}\phi_{2}\left(\frac{g}{k_{\mathrm{B}}T}\right),$ (7)
where $k_{\mathrm{B}}T$ is the usual thermal energy, the first three terms account for the ideal entropy of mixing and the last for the binding of mesogens and dopants to form dimers. The optimal fraction dimers $\phi_{2}$ minimizes the free energy Eq. 7 ,
$\frac{\phi_{2}}{\phi\hskip 2.84544pt\phi_{*}}=2\exp\left(\frac{g}{k_{\mathrm{B}}T}\right)\equiv 2K,$ (8)
where we defined the binding constant $K\equiv\exp(g/k_{\mathrm{B}}T)$.
It is useful to define the enantiomeric excess $\eta$, where we note that due to the presumed interconversion of P and M or $+$ and -- conformers, half the free mesogens are $+$ and half --. Let a chiral dopant bound to a mesogen turn this mesogen into a $+$ conformer. In that case $\phi_{+}=\frac{1}{2}\phi+\phi_{2}$ and $\phi_{-}=\frac{1}{2}\phi$, and $\eta=(\phi_{+}-\phi_{-})/\Phi=\phi_{2}/2\Phi$. Making use of the conservation of mass, we obtain from Eq. 8
$\frac{\eta}{(1-\eta)(\Phi_{*}-\eta\phi)}=K.$ (9)
This equation can be solved exactly,
$\eta=\frac{1}{2}(1+r+\varepsilon)-\frac{1}{2}\sqrt{(1+r+\varepsilon)^{2}-4r},$ (10)
where $r=\Phi_{*}/\Phi$ and $\varepsilon=1/K\Phi$. Notice that in the limit where the dopant concentration goes to zero, and $r=0$, we insist that the enantiomeric excess goes to zero too, $\eta=0$, in agreement with Eq. 10. In the limit of small dopant concentrations $r\ll 1$, we may Taylor expand Eq. 10 to give
$\eta\sim\frac{r}{1+\varepsilon}.$ (11)
The free energy gain of dimerization is equal to $-\frac{1}{2}g\phi_{2}=-g\eta\Phi$ per lattice site. This means that the free energy gain per mesogen must be equal to $-g\hskip 2.84544pt\eta$. In the limit $r\ll 1$, we can replace $\eta$ by Eq. 11. In the same limit $r=\Phi_{*}/\Phi=\Phi_{*}/(1-\Phi_{*})\sim\Phi*$ and $\varepsilon=1/K\Phi\sim 1/K$. If the binding is strong and $-g\gg k_{\mathrm{B}}T$, we have $\varepsilon\ll 1$, and the binding free energy per mesogen becomes equal to $-g\Phi_{*}$ which can still be much smaller than the thermal energy if $\Phi_{*}$ is sufficiently small. This is the expression that we use in the main text, where we realize that for the lattice model, the volume fraction dopant is equal to its mole fraction, $X=\Phi_{*}$.
Temperature dependence of the helical twisting power
According to our Maier-Saupe theory, the helical twisting power can be written as
$\beta=\frac{1}{1-\chi}\times\frac{q_{0}}{\pi}\times\left(\frac{g}{k_{\mathrm{B}}T}\right),$ (12)
where each term in the product should be expected to depend on the temperature $T$, since $\chi=\chi(T)$, $q_{0}=q_{0}(T)$ and $g=g(T)$. For instance, if we expand the last term around a reference temperature $T_{0}$ to linear order in $\Delta T/T_{0}=(T-T_{0})/T_{0}$, we get
$\frac{g(T)}{k_{\mathrm{B}}T}=\frac{g(T_{0})}{k_{\mathrm{B}}T_{0}}-\left(\frac{h(T_{0})}{k_{\mathrm{B}}T_{0}}\right)\frac{\Delta T}{T_{0}}+\cdots,$ (13)
where the free energy gain $g=h-Ts$ can be written in terms of the enthalpy gain $h$ and the entropy loss $s$ of binding. For binding to occur we must heave $g>0$. If $h>0$ and $s>0$ then the binding is enthalpy-driven, if $h<0$ and $s<0$ the binding is entropy-driven. If all other factors do not depend on temperature, then $\beta$ decreases with increasing temperature if the binding is entropy driven.
If we now consider the second term, and Taylor expand $q_{0}$ near $T=T_{0}$, we obtain
$q_{0}(T)=q_{0}(T_{0})+\frac{1}{J(T_{0})}\left(-\frac{1}{2}S_{K}(T_{0})+q_{0}(T_{0})S_{J}(T_{0}))\right)\frac{\Delta T}{T_{0}},$ (14)
where $K=H_{K}-TS_{K}$ and $J=H_{J}-TS_{J}$ in terms of the enthalpy gains $H_{J,K}(T)$ and entropy losses $S_{J,K}(T)$ associated with the chiral and nematic interaction strengths $K\geq 0$ and $J\geq 0$. If the interaction strengths $J$ and $K$ are purely enthalpic in nature and $S_{J,K}=0$, as is usually presumed, then $q_{0}$ does not depend on the temperature. This is not what is usually seen experimentally albeit that the temperature dependence can be quite weak over a large temperature range, and for some chiral nematogens $q_{0}$ may in fact switch sign at some temperature [39]. To account for any temperature dependence, the Maier-Saupe theory for cholesterics has to be extended [13]. We argue, however, that since $J$ and $K$ represent interactions in a coarse-grained model in which microscopic degrees of freedom have been tacitly “integrated out”, these quantities should be viewed as free energies rather than energies (or enthalpies). Hence, depending on the sign and magnitude of $S_{J}$ and $S_{K}$, $q_{0}$ may increase or decrease with increasing temperature. This is similar in nature to how the Flory-Huggins interaction parameter of polymer solutions acquires its temperature dependence, and may increase or decrease with increase temperature depending on solvent and polymer type, leading to upper and lower critical solution temperatures.
The temperature dependence of the first term is clearly non-trivial, at least in principle, because $\chi=q_{0}KS^{2}/k_{\mathrm{B}}T=2q_{0}^{2}S^{2}J/k_{\mathrm{B}}T$ is a function of a number of temperature dependent quantities. Even if $q_{0}$ does not appreciably depend on the temperature, the scalar nematic order parameter does, noting that $S=S(J/k_{\mathrm{B}}T)$ decreases with increasing temperature. Hence, we would expect $\chi$ to decrease with increasing temperature, and with that $(1-\chi)^{-1}$ too.
In conclusion, we expect $\beta$ to decrease with temperature, unless the binding of the dopant to the mesogens is entropy dominated and/or the magnitude of the pitch $q_{0}$ of the (hypothetical) enantiomerically pure compound increases with temperature, in which case $\beta$ may increase with temperature. In fact, we cannot exclude the possibility that depending on the specific temperature dependencies of $q_{0}$, $g$ and $\chi$, $\beta$ it may even vary non-monotonically with temperature as is sometimes seen experimentally [39].
Supplementary figures
Figure S1: Monte Carlo simulation results showing the enantiomeric excess, $\eta$, as a function of the dimensionless chiral biasing potential $g/k_{\mathrm{B}}T$. Configurations of $128^{3}$ sites were annealed from $T/T_{IN}=0.668$ to $T/T_{IN}=0.267$ for $2\times 10^{6}$ Monte Carlo steps at each temperature. Shown in solid black is the analytical function, $\eta=\tanh\frac{g}{k_{\mathrm{B}}T}$, representing the limit of zero chiral coupling, $K/J=0$.
Figure S2: Monte Carlo simulation results showing the coarsening behavior of chiral domains at a constant temperature and chiral interaction strength. Simulations of a box of $480^{3}$ sites were started in a racemic isotropic state, and run at a constant temperature of $T/T_{IN}=0.067$ and chiral interaction strength of $K/J=0.5$. Configurations were saved for visualization purposes every $1\times 10^{5}$ Monte Carlo steps.
Figure S3: Cropped cross-section along simulation box diagonal $(+1,-1,+1)$ of a Monte Carlo simulation of size $N=480$ with $K/J=0.5$ and $T/T_{IN}=0.067$ after $2\times 10^{6}$ Monte Carlo steps showing the director configuration in the deracemized cholesteric phase. The colors red and blue indicate the left- and right-handed enantiomeric state of the model mesogens.
Figure S4: Chiral susceptibility in the racemic nematic phase, $\chi_{c}$, for $K/J=0.5$ and critical transition temperature, $T_{\mathrm{c}}/T_{IN}=0.134$. Drawn lines show the mean-field exponent of $\gamma=1$ and fitted exponent of $\gamma=1.036$.
Figure S5: Monte Carlo simulation results showing the coarsening behavior of chiral domains at a constant temperature and chiral interaction strength. Simulations of size $480^{3}$ were started in the racemic isotropic state and run at a constant temperature of $T/T_{IN}=0.067$ and chiral interaction strength of $K/J=0.5$ with configurations saved for visualization every $1\times 10^{5}$ Monte Carlo steps.
Figure S6: Monte Carlo simulation results showing the final director, and chirality configurations at various values of the strength of the chiral interaction between nematogens. All simulations were annealed to a temperature of $T/T_{IN}=0.067$ from $T/T_{IN}=2.0$ in $30$ steps for $2\times 10^{6}$ Monte Carlo steps for a total of $6\times 10^{7}$ steps.
In this section we present additional material obtained from our simulations.
Fig. S1 presents additional plots of enantiomeric excess as a function of $g/k_{\mathrm{B}}T$. We include simulations with zero chiral coupling, $K/J=0.0$, to shown that as chiral interactions decrease, the enantiomeric excess approaches the analytical function, $\eta=\tanh\frac{g}{k_{\mathrm{B}}T}$. The enantiomeric excess was calculated from the last $6.25\%$ of each simulation’s annealing step.
Additional figures supporting the discussion of coarsening are shown in Fig. S2. We show the same, large, simulation of size $N=480$, but visualize the chiral domains through a cross-section, showing the $XZ$ plane through the middle of the simulation volume, with the colors blue and red representing the value of the Ising-like chirality variable describing left- and right-handedness of the model mesogens. We can see that, over time, chiral layers widen and the number of layers drop, that the layer normal is along the diagonal, and that edge dislocations between layers are recombining towards the open boundaries of the simulation volume.
Fig. S3 shows an inset of the $N=480$ simulation in its final configuration where the chiral layers are the widest. We visualize both the nematic director as a sphero-cylinder and the chiral state variable in color. The helical arrangement of the director field of the cholesteric is clearly shown, as is the sharp interface between the left- and right-handed cholesteric domains. We can also see that there are multiple pitches in each chiral domain, as well as a double-twist structure.
Fig. S4 shows a double-logarithmic plot of the chiral susceptibility, $\chi_{\eta}\sim|T-T_{\mathrm{c}}|^{-\gamma}$ as a function of the temperature $T-T_{\mathrm{c}}$ relative to the critical temperature $T_{\mathrm{c}}/T_{IN}=0.134$ in the racemic isotropic phase for the case $K/J=0.5$. The mean-field exponent of unity from fit of simulation data is shown in green and the fitted exponent is shown in blue.
Here, in Fig. S5, we continue to show the coarsening behavior of the large, $N=480$ simulation, visualized on the outer layer of chiral domains as shown in Fig. 4 c). The most prominent outer face shows that the chiral layer normal is along a body diagonal, while the energetically favorable axis is more complex along other faces such as the top. Layers may annihilate at the corners, due to the open boundary conditions allowing them to reduce their surface area easily here. This effect can especially be seen between Fig. S5 b-d).
Fig. S6 shows final director and chirality configurations for completely annealed $N=128$ systems in various locations of the phase diagram. Fig. S6a) is in the racemic nematic phase and, while it shows small clusters of like chirality, there is no long-range chiral ordering and the nematic director is clearly visible. Fig. S6 b) shows a system that has just transitioned into a deracemized cholesteric phase with the chiral domains encompassing multiple pitches. Fig. S6 c) shows a system that hints at transitioning to the deracemized cholesteric blue phase as shown in Fig. S6 d). All director and chirality configurations were vizualised as the horizontal cross-section through the middle of the simulation volume on the $XZ$ plane with director field colored with the value of the Ising-like chiral variable.